40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 27 18:03:47 2014, ericq, Update, Green Locking, uPDH Box Checkup uPDH_X.JPGuPDH_Y.JPGupdhTFs.pdfupdhNoises.pdfuPDH_Aug27.zip
    Reply  Thu Aug 28 17:34:20 2014, ericq, Update, Green Locking, uPDH Box Checkup TEK00006.PNG
       Reply  Fri Aug 29 17:28:07 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Fri Aug 29 23:53:35 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Tue Sep 2 20:02:06 2014, ericq, Update, Green Locking, uPDH Box Checkup 
       Reply  Wed Sep 3 00:17:22 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Wed Sep 3 20:38:40 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Fri Sep 5 04:07:44 2014, ericq, Update, Green Locking, X end uPDH Box Replaced 
                Reply  Fri Sep 5 16:13:00 2014, ericq, Update, Green Locking, new X Green PDH measurements 6x
                   Reply  Sat Sep 6 01:48:54 2014, rana, Update, Green Locking, new X Green PDH measurements 
                      Reply  Sat Sep 6 14:49:12 2014, ericq, Update, Green Locking, new X Green PDH measurements xGComparison.pdfxGspectra.pdf
                         Reply  Sun Sep 7 17:06:38 2014, rana, Update, Green Locking, new X Green PDH measurements 
                         Reply  Mon Sep 8 14:14:13 2014, Jenne, Update, Green Locking, new Y Green PDH measurements 6x
Message ID: 10457     Entry time: Fri Sep 5 04:07:44 2014     In reply to: 10447     Reply to this: 10460
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: X end uPDH Box Replaced 

Just a quick note, plots and data will come tomorrow:

I grabbed an unused uPDH board from the ATF (thanks Zach!), and re-stuffed almost the entire thing to match Jenne's latest schematic for the y end box. I also threw some 22uF caps on the regulators, as Koji did with the previous box, to eliminate some oscillations up in the high 10s of kHz. I replaced the tragedy of a box that I created on Wednesday with this new box. The arm locks pretty stably with the boost on, 30 degrees of phase margin with 10kHz UGF, and locks pretty darn reliably. 

Now we should now have two nicely boosted PDH loops. I'll do a noise/loop breakdown again in the upcoming days. 

ELOG V3.1.3-